Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability doubt

Expert replies
by sachin_yadav » Wed Dec 08, 2010 7:36 pm
Hi All,

Came across a question from "Veritas prep".

A telephone number contains 10 digits, including a 3-digit area code. Bob remembers the area code and the next 5 digits of the number. He also remembers that the remaining digits are not 0, 1, 2, 5, or 7. If bob tries to find the number by guessing the remaining digits at random, the probability that he will be able to find the correct number in at most 2 attempts is closest to which of the following ?

(A) 1/625
(B) 2/625
(C) 4/625
(D) 25/625
(E) 50/625


Thanks
Sachin
Join the discussion
Source: — Problem Solving |

by shovan85 » Wed Dec 08, 2010 8:10 pm
Numbers left to fill last two digits are 3,4,6,8,9.
So, Total 5 numbers are left to be chosen.

Probability of choosing right numbers in two places
= Probability of choosing right number in first place * Probability of choosing right number in second place
=1/5 * 1/5
= 1/25

Thus, Probability of not choosing right numbers in two places = 1-1/25 = 24/25

Two attempts and it has to succeed thus

1) 1st Attempt: Wrong, 2nd Attempt: Correct
So, P = 24/25 * 1/25 = 24/625

2) 1st Attempt: Correct
So, P = 1/25

Thus requiredd probability 24/625 + 1/25 = 49/625

IMO E Nearest One
If the problem is Easy Respect it, if the problem is tough Attack it
Join the discussion

by beat_gmat_09 » Wed Dec 08, 2010 8:24 pm
Numbers to fill - 3,4,6,8,9
5 can be filled in 5*5 = 25 ways, assuming repetition.

Probability of getting in right in 2nd attempt is just as selecting 1 option from 25 Proba = 1/25
Probab of getting right in 1st attempt = 1/25

Total probab = 1/25 + 1/25 = 2/25
2/25 *25/25 = 50/625
IMO E
Hope is the dream of a man awake
Join the discussion

by Night reader » Wed Dec 08, 2010 8:45 pm
sachin_yadav wrote:Hi All,

Came across a question from "Veritas prep".

A telephone number contains 10 digits, including a 3-digit area code. Bob remembers the area code and the next 5 digits of the number. He also remembers that the remaining digits are not 0, 1, 2, 5, or 7. If bob tries to find the number by guessing the remaining digits at random, the probability that he will be able to find the correct number in at most 2 attempts is closest to which of the following ?

(A) 1/625
(B) 2/625
(C) 4/625
(D) 25/625
(E) 50/625


Thanks
Sachin
5 digits are left =>
there are 5 ways to arrange 5 digits for each of two units (last) in the number

___ ___ => 5*5 => 25
P(1st attempt) = 1/25

P(1st attempt wrong) = 1-1/25 = 24/25

P(2nd attempt) = 1/24

max P(correct @2nd attempt) = (24/25)*(1/24) =24/600

Combining two probabilities 24/600 + 1/25 = (24+24)/600 = 48/600 = 12.5

50/625 = 48/600

Answer E is correct
Join the discussion

by sachin_yadav » Wed Dec 08, 2010 10:36 pm
Thank you guys. I really appreciate for the help, but probability is definitely a nightmare for me.

Regards,
Sachin
Join the discussion